Solve the system by addition. 5x – 8y = –14 x – 2y = –6
Answer by checkley71(8403)

5X-8Y=-14
X-2Y=-6 THEN X=2Y-6 AND WE SUBSTITUTE (2Y-6) FOR X IN THE OTHER EQUATION
WE GET
5(2Y-6)-8Y=-14
10Y-30-8Y=-14
10Y-8Y=30-14
2Y=16
Y=16/2
Y=8 SOLUTION
X=2*8-6
X=16-6
X=10 SOLUTION
PROOF
5*10-8*8=-14
50-64=-14
-14=-14